OpenStudy (angelab97):

Josh goes on a date with Lisette on the first day, Angelique on the second day, Fantasia on the third day, Raquel on the fourth day, Patrice on the fifth day, and Shayla on the sixth day. He begins the patterns again on the seventh day, and repeats it over and over again until he gets hit by a bus. If he was on his way home from a date with Fantasia when he got hit which of the following could be the number dates he went on before his accident?

OpenStudy (angelab97):

@undeadknight26

OpenStudy (angelab97):

is there anyway i can solve this algebraically and easily?

undeadknight26 (undeadknight26):

This is a confusing question because a bus bus shouldn;t hit a guy so much...@AngelaB97 is this a serious question?

OpenStudy (angelab97):

yes it is a SAT question and if you can't answer it, it's fine thanks anyways

OpenStudy (anonymous):

i think it should be 2 O_o

OpenStudy (angelab97):

the answer choices are 19,23,84,173,279

undeadknight26 (undeadknight26):

Well without knowing how many times exactly it repeats its practically impossible... If anything its 84 or 279 84/3 = 28 279/3 = 93

undeadknight26 (undeadknight26):

3 since fantiza is the 3rd day.

OpenStudy (angelab97):

the answer is supposed to be 279 i just don't know how to get there

OpenStudy (mathstudent55):

Are you given choices?

OpenStudy (angelab97):

19, 23, 84, 173, 29 @mathstudent55

OpenStudy (angelab97):

19,23,84,173,279 @mathstudent55

OpenStudy (mathstudent55):

Day Person 1 Lisette 2 Angelique 3 Fantasia 4 Raquel 5 Patrice 6 Shayla 7 Lisette 8 Angelique 9 Fantasia 10 Raquel 11 Patrice 12 Shayla etc.

OpenStudy (mathstudent55):

Fantasia is on day 3, 9, 15, 21, ... 19 mod 6 = 1 23 mod 6 = 5 84 mod 6 = 0 173 mod 6 = 5 279 mod 6 = 3 The only one whose remainder is 3 is 279.

undeadknight26 (undeadknight26):

I was close but not close enough.

OpenStudy (mathstudent55):

The answer has to be a number that when divided by 6 gives you a remainder of 3.

OpenStudy (the_fizicx99):

Simple pattern, Day1 Lisette , Day 2Angelique, \(\ Day ~ \color{red}{3} ~ Fantasia\), Day 4 Raquel, Day 5 Patrice, Day 6 Shayla and it starts over. Day 7 Lisette , Day 8 Angelique, \(\ Day ~ \color{red}{9} ~ Fantasia\), Day 10 Raquel, Day 11 Patrice, Day 12 Shayla. Fantasia's repeats every 6th day, although it will always be on a multiple of 3, it won't be every multiple of 3. If that makes sense to you. Because it's -3 it has to be a multiple of 6, 19 - 3 = 16/3 Won't work 23 - 3 - 20/3 Won't work 84 - 3 = 81/3 won't work 279 - 3 = 276/3 will work

OpenStudy (angelab97):

alrigt, thanks for everyone who answered!! @mathstudent55 @tHe_FiZiCx99 @Chris01 @undeadknight26

OpenStudy (mathstudent55):

You're welcome.